1. (This problem can be done using either a spreadsheet or your financial calculator. In either case, show

all of your work. In the case of a spreadsheet, please print out the entire spreadsheet and indicate

where your answers to the various parts of this problem can be found.)

A CPM loan is made for $50,000 at 5 percent interest for 30 years. Calculate:

a. The monthly payments for principal and interest.

b. Interest and principal payments during month 1.

c. The outstanding loan balance if the loan is repaid at the end of year 6.

d. What would the breakdown of interest and principal be during month 25?
Answer #1

Solution:

(a) B8 XPMT(B4/B6,B5 B6,-B3) 3 Loan amount 4 Intereset rate 5 Maturity period in yrs $50,000 5% 30 Compounding and 6 payments in a year 12 8 (a)Monthly payments $268.41

(b) Interest payment during month 1 = ($50,000*5%)*1/12 = $208.33

   Principal payment = $268.41 - 208.33 = $60.08

(c) The outstanding balance using prospective method = Discounted value of remaining payments

   = $268.41[1-(1+0.05/12)-(30-6)*12] / (0.05/12)

   = $44,967.65

(d) Interest in the 25th pmt =IPMT(5%/12,25,360,-50000) (using excel )

   = $202.03

Principal = PPMT(5%/12,25,360,-50000)

= $66.38